User Tools

Site Tools


lbaops:lbajun2016:v493cpalog

Parkes had wrong patching for approx first 1.5 hrs

Parkes Pols are swapped

UT 05:15

Generator failure alarm. Parkes Telescope went into stow position.

UT 05:30

Parkes Telescope unstowed and slewing to source.

at 18:24 UT noticed Pictor VNC had a flashing message saying “bank/vsn ‘a/pks+0015’ should be changed”, with buttons to “continue” or “stop flashing”. I chose continue assuming the msg was advisory, and there were no subsequent flashing messages or apparent error messages in the logs. Await advice on whether disk actually filled and data no longer recorded, or whether ok… Not clear how long that message had been flashing (that part of screen was hidden under web browser)

lbaops/lbajun2016/v493cpalog.txt · Last modified: 2016/06/30 07:23 by pge